A Carnot engine operating between 527C and 27 C produces 600KW of work. Determine,

(a) Qa, KW; (b) Qr, KW; (c) ΔS during heat rejection, KW/K and (d) thermal efficiency of the cycle.


1
Expert's answer
2022-04-04T09:03:55-0400

"\eta=1-\frac {T_2}{T_1}=0.625,"

"Q_a=600~kW,"

"Q_r=\frac {Q_a}{1-\eta}=1600~kW,"

"∆S=\frac{Q_r}{∆T}=3.2~\frac WK."